OpenStudy (mathmath333):

How to quickly calculate 5!/e

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ How to quickly calculate}\hspace{.33em}\\~\\ & \lfloor{\dfrac{5!}{e}\rfloor} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

by hand

OpenStudy (mathmate):

Do you mean pen and paper or mentally?

OpenStudy (mathmath333):

by pen and paper (hand)

OpenStudy (mathmate):

So would that not be 120/2.718281828?

OpenStudy (mathmath333):

yes

OpenStudy (mathmate):

Or you can do a quick mental approximation by 120/2.7=40/0.9=44.4

OpenStudy (mathmate):

If you want to refine it, subtract 2/300 of the calculated value (mentally if you want) to get 44.15 Accurate value is 44.14553294057308

OpenStudy (mathmath333):

it should be correct

OpenStudy (mathmate):

Interesting problem. Will be back.

OpenStudy (mathmath333):

ur thinking is correct of \(\approx n/2.7\)

OpenStudy (anonymous):

trying to learn derangements? :)

OpenStudy (mathmate):

As long as it is between 44 and 45, the floor function will convert it to 44. We need additional accuracy to get the lower bound so it doesn't go below 44.

OpenStudy (mathmath333):

nice forecasting of dearrangments.